OpenStudy (anonymous):

A quadratic equation in form ax^2+bx+c=0 cannot have: one real solution one imaginary solution two real solutions two imaginary solutions

OpenStudy (anonymous):

a and b are answers

OpenStudy (anonymous):

which ones srry I didn't label them

OpenStudy (anonymous):

one real //////////one imginary

OpenStudy (anonymous):

its one or the other

OpenStudy (anonymous):

It can have one real solution if you ignore "repeated" roots.

OpenStudy (anonymous):

so one inaginary

OpenStudy (anonymous):

Yup, if you twist my arm.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

That was correct thanks

OpenStudy (anonymous):

I do have a certain amount of sympathy with html1618's position.

OpenStudy (anonymous):

It cannot have 1 imaginary solution as imaginary solutions come from cubics. \[ax^3 + bx^2 + cx + d\]

OpenStudy (anonymous):

@Twill What are the solutions of x^2 +2?

OpenStudy (anonymous):

2 of 'em, right?

OpenStudy (anonymous):

@estudier i >.>

OpenStudy (anonymous):

well, x^2 + 2 = 0 x^2 = -2 So x = i

OpenStudy (anonymous):

Well, not quite; try using the quadratic formula, see what you get...

OpenStudy (anonymous):

+/- 2!

OpenStudy (anonymous):

I thought there was a square root sign in the quadratic formula?

OpenStudy (anonymous):

Sorry, Pie_Boi, didn't mean to hijack your question.

OpenStudy (anonymous):

\[-b +/- \sqrt{b^2 - 4ac}\div2a\] a =1 b = 0 c = 2 \[\pm \sqrt{(0)^2 - 4(1)(2)} \div2(1)\] \[\pm \sqrt{-8}\div2\] \[\pm \sqrt{-4}\] Gives you i?

OpenStudy (anonymous):

I can write \[\sqrt{-4}\] as \[\sqrt{-1}\sqrt{4}\] Do you agree?

OpenStudy (anonymous):

Yes, fair point. So i and \[\pm2?\]

OpenStudy (anonymous):

Usually written as \[\pm i \sqrt{2}\]

OpenStudy (anonymous):

Well done.

OpenStudy (anonymous):

Thank you! Sorry for the Hijack Pie_Boi

OpenStudy (anonymous):

no problem at all :)